sqrt of x-2=1
Answer:
Show work:

Answer by ankor@dixie-net.com(12692)

sqrt of x-2=1
:
Really this is very easy:
sqrt(x-2) = 1
:
Squaring both sides:
(sqrt(x-2))^2 = 1^2
:
That gets rid of the radical and you have:
x - 2 = 1
:
Add 2 to both sides:
x = 1 + 2
x = 3
:
Check in the original equation, substitute 3 for x:
sqrt(3-2) = 1
sqrt(1) = 1, obviously